# Remove that Square Root: A New Efficient Scale-Invariant Version of AdaGrad

### v1

@@ 1 Introduction
where $f:\mathbb{R}^{d}\to\mathbb{R}$ is a $L$-smooth and generally non-convex function. In particular, we are interested in the situations when the objective has either expectation $f(w)=\mathbb{E}_{\xi\sim{\cal D}}[f_{\xi}(w)]$ or finite-sum $f(w)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(w)$ form. Such minimization problems are crucial in machine learning, where $w$ corresponds to the model parameters. Solving these problems with stochastic gradient-based optimizers has gained much interest owing to their wider applicability and low computational cost. Stochastic Gradient Descent (SGD[Robbins and Monro, 1951] and similar algorithms require the knowledge of parameters like $L$ for convergence and are very sensitive to the choice of the stepsize in general. Therefore, SGD requires hyperparameter tuning, which can be computationally expensive. To address these issues, it is common practice to use adaptive variants of stochastic gradient-based methods that can converge without knowing the function’s structure.
There exist many adaptive algorithms such as AdaGrad [Duchi et al., 2011], Adam [Kingma and Ba, 2014], AMSGrad [Reddi et al., 2019], D-Adaptation [Defazio and Mishchenko, 2023], Prodigy [Mishchenko and Defazio, 2023] and their variants. These adaptive techniques are capable of updating their step sizes on the fly. For instance, the AdaGrad method determines its step sizes using a cumulative sum of the coordinate-wise squared (stochastic) gradient of all the previous iterates:
$\text{{\color[rgb]{0,110,51}\definecolor[named]{pgfstrokecolor}{rgb}{0,110,51}\pgfsys@color@rgb@stroke{0}{110}{51}\pgfsys@color@rgb@fill{0}{110}{51}\sf{\color[rgb]{0,110,51}\definecolor[named]{pgfstrokecolor}{rgb}{0,110,51}\pgfsys@color@rgb@stroke{0}{110}{51}\pgfsys@color@rgb@fill{0}{110}{51}AdaGrad}}:}\;w_{t+1}=w_{t}-\frac{\beta g_{t}}{\sqrt{\text{ diag}\left(\Delta I+\sum\limits_{\tau=1}^{t}g_{\tau}g_{\tau}^{\top}\right)}},$ (2)
where $g_{t}$ represents an unbiased estimator of $\nabla f(w_{t})$, i.e., $\mathbb{E}\left[g_{t}\mid w_{t}\right]=\nabla f(w_{t})$, $\text{diag}(M)\in\mathbb{R}^{d}$ is a vector of diagonal elements of matrix $M\in\mathbb{R}^{d\times d}$, $\Delta>0$, and the division by vector is done component-wise. Ward et al. [2020] has shown that this method achieves a convergence rate of $\mathcal{O}\left(\nicefrac{{\log T}}{{\sqrt{T}}}\right)$ for smooth functions, similar to SGD, without prior knowledge of the functions’ parameters. However, the performance of AdaGrad deteriorates when applied to data that may exhibit poor scaling or ill-conditioning. In this work, we propose a novel algorithm, KATE, to address the issues of poor data scaling. KATE is also a stochastic adaptive algorithm that can achieve a convergence rate of $\mathcal{O}\left(\nicefrac{{\log T}}{{\sqrt{T}}}\right)$ for smooth non-convex functions in terms of $\min_{t\in[T]}\mathbb{E}\left[\left\|\nabla f(w_{t})\right\|\right]^{2}$.
@@ 1.1 Related Work
A significant amount of research has been done on adaptive methods over the years, including AdaGrad [Duchi et al., 2011, McMahan and Streeter, 2010], AMSGrad [Reddi et al., 2019], RMSProp [Tieleman and Hinton, 2012], AI-SARAH [Shi et al., 2023], and Adam [Kingma and Ba, 2014]. However, all these works assume that the optimization problem is contained in a bounded set. To address this issue, Li and Orabona [2019] proposes a variant of the AdaGrad algorithm, which does not use the gradient of the last iterate (this makes the step sizes of $t$-th iteration conditionally independent of $g_{t}$) for computing the step sizes and proves convergence for the unbounded domain.
Each of these works considers a vector of step sizes for each coefficient. Duchi et al. [2011] and McMahan and Streeter [2010] simultaneously proposed the original AdaGrad algorithm. However, McMahan and Streeter [2010] was the first to consider the vanilla scalar form of AdaGrad, known as
$\text{{\color[rgb]{0,110,51}\definecolor[named]{pgfstrokecolor}{rgb}{0,110,51}\pgfsys@color@rgb@stroke{0}{110}{51}\pgfsys@color@rgb@fill{0}{110}{51}\sf{\color[rgb]{0,110,51}\definecolor[named]{pgfstrokecolor}{rgb}{0,110,51}\pgfsys@color@rgb@stroke{0}{110}{51}\pgfsys@color@rgb@fill{0}{110}{51}AdaGradNorm}}: }w_{t+1}=w_{t}-\frac{\beta g_{t}}{\sqrt{\Delta+\sum_{\tau=0}^{t}\left\|g_{\tau}\right\|^{2}}}.$ (3)
@@
Recently, Defazio and Mishchenko [2023] introduced the D-Adaptation method, which has gathered considerable attention due to its promising empirical performances. In order to choose the adaptive step size optimally, one requires knowledge of the initial distance from the solution, i.e., $D\coloneqq\left\|w_{0}-w_{*}\right\|$ where $w_{*}\in\text{argmin}_{w\in\mathbb{R}^{d}}f(w)$. The D-Adaptation method works by maintaining an estimate of $D$ and the stepsize choice in this case is $\nicefrac{{d_{t}}}{{\sqrt{\sum_{\tau=0}^{t}\|g_{\tau}\|^{2}}}}$ for the $t$-th iteration. [^3] Here $d_{t}$ is an estimate of $D$. Mishchenko and Defazio [2023] further modifies the algorithm in a follow-up work and introduces Prodigy (with stepsize choice $\nicefrac{{d^{2}_{t}}}{{\sqrt{\sum_{\tau=0}^{t}d_{\tau}^{2}\|g_{\tau}\|^{2}}}}$) to improve the convergence speed.
Table 1: Summary of convergence guarantees for closely-related adaptive algorithms to solve smooth non-convex stochastic optimization problems. Convergence rates are given in terms of $\min_{t\in[T]}\mathbb{E}\left[\left\|\nabla f(w_{t})\right\|\right]^{2}$. Moreover, we highlight our method’s scale-invariance property.
[htb] Algorithm Convergence rate Scale invariance AdaGradNorm [Ward et al., 2020] $\mathcal{O}\left(\nicefrac{{\log T}}{{\sqrt{T}}}\right)$ AdaGrad [Défossez et al., 2020] $\mathcal{O}\left(\nicefrac{{\log T}}{{\sqrt{T}}}\right)$ Adam [Défossez et al., 2020] $\mathcal{O}\left(\nicefrac{{\log T}}{{\sqrt{T}}}\right)$ KATE (this work) $\mathcal{O}\left(\nicefrac{{\log T}}{{\sqrt{T}}}\right)$
Another exciting line of work on adaptive methods is Polyak stepsizes. Polyak [1969] first proposed Polyak stepsizes for subgradient methods, and recently, the stochastic version (also known as SPS) was introduced by Oberman and Prazeres [2019], Loizou et al. [2021] and Gower et al. [2021]. For a finite sum problem $\min_{w\in\mathbb{R}^{d}}f(w):=\frac{1}{n}\sum_{i=1}^{n}f_{i}(w)$, Loizou et al. [2021] uses $\frac{f_{i}(w_{k})-f_{i}^{*}}{c\|\nabla f_{i}(w^{k})\|^{2}}$ as their stepsize choices, [^4] Here $f_{i}^{*}:=\min_{w\in\mathbb{R}^{d}}f_{i}(w)\quad\forall i\in[n]$. while Oberman and Prazeres [2019] uses $\frac{2(f(w_{k})-f^{*})}{\mathbb{E}\left[\|\nabla f_{i}(w^{k})\|^{2}\right]}$ for $k$-th iteration. However, these methods are impractical when $f^{*}$ or $f_{i}^{*}$ is unknown. Following its introduction, several variants of the SPS algorithm emerged [Li et al., 2022, D’Orazio et al., 2021]. Lately, Orvieto et al. [2022] tackled the issues with unknown $f_{i}^{*}$ and developed a truly adaptive variant. In practice, the SPS method shows excellent empirical performance on overparameterized deep learning models (which satisfy the interpolation condition i.e. $f_{i}^{*}=0\quad\forall i\in[n]$) [Loizou et al., 2021].
@@ 1.2 Main Contribution
We prove that for smooth non-convex problems with noise having bounded variance KATE has ${\cal O}(\nicefrac{{\log(T)}}{{\sqrt{T}}})$ convergence rate (Theorem 3.4), matching the best-known rates for AdaGrad and Adam [Défossez et al., 2020].
@@ 2 Motivation and Algorithm Design
Here, we consider the parameter estimation problem in generalized linear models (GLMs) [Nelder and Wedderburn, 1972, Agresti, 2015] using maximum likelihood estimation. GLMs are an extension of linear models and encompass several other valuable models, such as logistic [Hosmer Jr et al., 2013] and Poisson regression [Frome, 1983], as special cases. The parameter estimation to fit GLM on dataset $\{x_{i},y_{i}\}_{i=1}^{n}$ (where $x_{i}\in\mathbb{R}^{d}$ are feature vectors and $y_{i}$ are response variables) can be reformulated as
@@
for differentiable functions $\varphi_{i}:\mathbb{R}\to\mathbb{R}$ [Shalev-Shwartz and Ben-David, 2014]. For example, the linear regression on data $\{x_{i},y_{i}\}_{i=1}^{n}$ is equivalent to solving (4) with $\varphi_{i}(z)=(z-y_{i})^{2}$. Next, the choice of $\varphi_{i}$ for logistic regression is $\varphi_{i}(z)=\log\left(1+\exp{\left(-y_{i}z\right)}\right)$.
@@
respectively, for functions $\varphi_{i}$. In this work, we want to design an algorithm with equivalent performance for the problems (4) and (5). If we can do that, the new algorithm’s performance will not deteriorate for poorly scaled data. To develop such an algorithm, we replace the denominator of AdaGrad step size with its square (remove the square root from the denominator), i.e., $\forall k\in[d]$
$w_{t+1}[k]$ $=$ $w_{t}[k]-\frac{\beta m_{t}[k]}{\sum_{\tau=0}^{t}g_{\tau}^{2}[k]}g_{t}[k].$ (6)
for some [^5] Sequence $\{m_{t}\}_{t\geq 0}$ can depend on the problem but is assumed to be scale-invariant. $m_{t}\in\mathbb{R}^{d}$. The following proposition shows that this method (6) satisfies a scale-invariance property with respect to functional value.
@@
$\hat{w}_{t+1}[k]$ $=$ $\hat{w}_{t}[k]-\frac{\beta m_{t}[k]}{\sum_{\tau=0}^{t}g_{\tau}^{2}[k]}g_{t}[k],$ (7)
$\quad\hat{w}^{V}_{t+1}[k]$ $=$ $\hat{w}^{V}_{t}[k]-\frac{\beta m_{t}[k]}{\sum_{\tau=0}^{t}\left(g_{\tau}^{V}[k]\right)^{2}}g^{V}_{t}[k]$ (8)
@@
$\hat{w}_{t}=V\hat{w}^{V}_{t},\quad Vg_{t}=g^{V}_{t},\quad f\left(\hat{w}_{t}\right)=f^{V}\left(\hat{w}^{V}_{t}\right)$
@@
Proof.
$\hat{w}_{1}[k]$ $=$ $\frac{-\beta m_{0}[k]\varphi^{\prime}_{i_{0}}(0)x_{i_{0}}[k]}{\left(\varphi^{\prime}_{i_{0}}(0)x_{i_{0}}[k]\right)^{2}}=\frac{-\beta m_{0}[k]}{\varphi^{\prime}_{i_{0}}(0)x_{i_{0}}[k]},$
$\hat{w}^{V}_{1}[k]$ $=$ $\frac{-\beta m_{0}[k]\varphi^{\prime}_{i_{0}}(0)V_{kk}x_{i_{0}}[k]}{\left(\varphi^{\prime}_{i_{0}}(0)V_{kk}x_{i_{0}}[k]\right)^{2}}=\frac{-\beta m_{0}[k]}{\varphi^{\prime}_{i_{0}}(0)V_{kk}x_{i_{0}}[k]}.$
$Vg_{1}$ $=$ $\varphi^{\prime}_{i_{1}}\left(x_{i_{1}}^{\top}\hat{w}_{1}\right)Vx_{i_{1}}$ (10)
$=$ $\varphi^{\prime}_{i_{1}}\left(x_{i_{1}}^{\top}V\hat{w}^{V}_{1}\right)Vx_{i_{1}}$
$=$ $g_{1}^{V},$
$\hat{w}_{t+1}[k]$ $=$ $\hat{w}_{t}[k]-\frac{\beta m_{t}[k]}{\sum_{\tau=0}^{t}g^{2}_{\tau}[k]}g_{t}[k]$
$=$ $V_{kk}\hat{w}^{V}_{t}[k]-\frac{\beta m_{t}[k]V_{kk}^{2}}{\sum_{\tau=0}^{t}\left(g^{V}_{\tau}[k]\right)^{2}}\frac{g^{V}_{t}[k]}{V_{kk}}$
$=$ $V_{kk}\hat{w}^{V}_{t+1}[k].$
$f(\hat{w}_{t})$ $=$ $\frac{1}{n}\sum_{i=1}^{n}\varphi_{i}\left(x_{i}^{\top}\hat{w}_{t}\right)$
$=$ $\frac{1}{n}\sum_{i=1}^{n}\varphi_{i}\left(x_{i}^{\top}V\hat{w}^{V}_{t}\right)$
$=$ $f^{V}\left(\hat{w}^{V}_{t}\right).$
$\left\|g_{t}^{V}\right\|_{V^{-2}}^{2}$ $=$ $\left(g_{t}^{V}\right)^{\top}V^{-2}g_{t}^{V}$
$=$ $g_{t}^{\top}VV^{-2}Vg_{t}$
$=$ $\left\|g_{t}\right\|^{2}.$
This completes the proof of Proposition 2.1. ∎
The Proposition 2.1 highlights that the update rule of the form (6) satisfies a scale-invariance property for GLMs.

### v4

@@ 1 Introduction
where $f:\mathbb{R}^{d}\to\mathbb{R}$ is a $L$-smooth and generally non-convex function. In particular, we are interested in the situations when the objective has either expectation $f(w)=\mathbb{E}_{\xi\sim{\cal D}}[f_{\xi}(w)]$ or finite-sum $f(w)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(w)$ form. Such minimization problems are crucial in machine learning, where $w$ corresponds to the model parameters. Solving these problems with stochastic gradient-based optimizers has gained much interest owing to their wider applicability and low computational cost. Stochastic Gradient Descent (SGD(Robbins and Monro, , 1951) and similar algorithms require the knowledge of parameters like $L$ for convergence and are very sensitive to the choice of the stepsize in general. Therefore, SGD requires hyperparameter tuning, which can be computationally expensive. To address these issues, it is common practice to use adaptive variants of stochastic gradient-based methods that can converge without knowing the function’s structure.
There exist many adaptive algorithms such as AdaGrad (Duchi et al., , 2011), Adam (Kingma and Ba, , 2014), AMSGrad (Reddi et al., , 2019), D-Adaptation (Defazio and Mishchenko, , 2023), Prodigy (Mishchenko and Defazio, , 2023), AI-SARAH (Shi et al., , 2023) and their variants. These adaptive techniques are capable of updating their step sizes on the fly. For instance, the AdaGrad method determines its step sizes using a cumulative sum of the coordinate-wise squared (stochastic) gradient of all the previous iterates:
$\text{{\color[rgb]{0,0.4296875,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4296875,0.19921875}\sf{\color[rgb]{0,0.4296875,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4296875,0.19921875}AdaGrad}}:}\;w_{t+1}=w_{t}-\frac{\beta g_{t}}{\sqrt{\text{ diag}\left(\Delta I+\sum\limits_{\tau=1}^{t}g_{\tau}g_{\tau}^{\top}\right)}},$ (2)
where $g_{t}$ represents an unbiased estimator of $\nabla f(w_{t})$, i.e., $\mathbb{E}\left[g_{t}\mid w_{t}\right]=\nabla f(w_{t})$, $\text{diag}(M)\in\mathbb{R}^{d}$ is a vector of diagonal elements of matrix $M\in\mathbb{R}^{d\times d}$, $\Delta>0$, and the division by vector is done component-wise. Ward et al., (2020) has shown that this method achieves a convergence rate of $\mathcal{O}\left(\nicefrac{{\log T}}{{\sqrt{T}}}\right)$ for smooth functions, similar to SGD, without prior knowledge of the functions’ parameters. However, the performance of AdaGrad deteriorates when applied to data that may exhibit poor scaling or ill-conditioning. In this work, we propose a novel algorithm, KATE, to address the issues of poor data scaling. KATE is also a stochastic adaptive algorithm that can achieve a convergence rate of $\mathcal{O}\left(\nicefrac{{\log T}}{{\sqrt{T}}}\right)$ for smooth non-convex functions in terms of $\min_{t\in[T]}\mathbb{E}\left[\left\|\nabla f(w_{t})\right\|\right]^{2}$.
@@ 1.1 Related Work
A significant amount of research has been done on adaptive methods over the years, including AdaGrad (Duchi et al., , 2011; McMahan and Streeter, , 2010), AMSGrad (Reddi et al., , 2019), RMSProp (Tieleman and Hinton, , 2012), AI-SARAH (Shi et al., , 2023), and Adam (Kingma and Ba, , 2014). However, all these works assume that the optimization problem is contained in a bounded set. To address this issue, Li and Orabona, (2019) proposes a variant of the AdaGrad algorithm, which does not use the gradient of the last iterate (this makes the step sizes of $t$-th iteration conditionally independent of $g_{t}$) for computing the step sizes and proves convergence for the unbounded domain.
Each of these works considers a vector of step sizes for each coefficient. Duchi et al., (2011) and McMahan and Streeter, (2010) simultaneously proposed the original AdaGrad algorithm. However, McMahan and Streeter, (2010) was the first to consider the vanilla scalar form of AdaGrad, known as
$\text{{\color[rgb]{0,0.4296875,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4296875,0.19921875}\sf{\color[rgb]{0,0.4296875,0.19921875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.4296875,0.19921875}AdaGradNorm}}: }w_{t+1}=w_{t}-\frac{\beta g_{t}}{\sqrt{\Delta+\sum_{\tau=0}^{t}\left\|g_{\tau}\right\|^{2}}}.$ (3)
@@
Recently, Defazio and Mishchenko, (2023) introduced the D-Adaptation method, which has gathered considerable attention due to its promising empirical performances. In order to choose the adaptive step size optimally, one requires knowledge of the initial distance from the solution, i.e., $D\coloneqq\left\|w_{0}-w_{*}\right\|$ where $w_{*}\in\text{argmin}_{w\in\mathbb{R}^{d}}f(w)$. The D-Adaptation method works by maintaining an estimate of $D$ and the stepsize choice in this case is $\nicefrac{{d_{t}}}{{\sqrt{\sum_{\tau=0}^{t}\|g_{\tau}\|^{2}}}}$ for the $t$-th iteration (here $d_{t}$ is an estimate of $D$). Mishchenko and Defazio, (2023) further modifies the algorithm in a follow-up work and introduces Prodigy (with stepsize choice $\nicefrac{{d^{2}_{t}}}{{\sqrt{\sum_{\tau=0}^{t}d_{\tau}^{2}\|g_{\tau}\|^{2}}}}$) to improve the convergence speed.
Table 1: Summary of convergence guarantees for closely-related adaptive algorithms to solve smooth non-convex stochastic optimization problems. Convergence rates are given in terms of $\min_{t\in[T]}\mathbb{E}\left[\left\|\nabla f(w_{t})\right\|\right]^{2}$. We highlight KATE’s scale-invariance property for problems of type (4).
Algorithm Convergence rate Scale invariance
AdaGradNorm (Ward et al., , 2020) $\mathcal{O}\left(\nicefrac{{\log T}}{{\sqrt{T}}}\right)$
AdaGrad (Défossez et al., , 2020) $\mathcal{O}\left(\nicefrac{{\log T}}{{\sqrt{T}}}\right)$
Adam (Défossez et al., , 2020)
$\mathcal{O}\left(\nicefrac{{\log T}}{{\sqrt{T}}}\right)$
KATE (this work) $\mathcal{O}\left(\nicefrac{{\log T}}{{\sqrt{T}}}\right)$
Another exciting line of work on adaptive methods is Polyak stepsizes. Polyak, (1969) first proposed Polyak stepsizes for subgradient methods, and recently, the stochastic version (also known as SPS) was introduced by Oberman and Prazeres, (2019); Loizou et al., (2021); Abdukhakimov et al., (2024, 2023); Li et al., (2023) and Gower et al., (2021). For a finite sum problem $\min_{w\in\mathbb{R}^{d}}f(w):=\frac{1}{n}\sum_{i=1}^{n}f_{i}(w)$, Loizou et al., (2021) uses $\frac{f_{i}(w_{t})-f_{i}^{*}}{c\|\nabla f_{i}(w_{t})\|^{2}}$ as their stepsize choices (here $f_{i}^{*}:=\min_{w\in\mathbb{R}^{d}}f_{i}(w)$), while Oberman and Prazeres, (2019) uses $\frac{2(f(w_{t})-f^{*})}{\mathbb{E}\left[\|\nabla f_{i}(w_{t})\|^{2}\right]}$ for $k$-th iteration. However, these methods are impractical when $f^{*}$ or $f_{i}^{*}$ is unknown. Following its introduction, several variants of the SPS algorithm emerged (Li et al., , 2023; D’Orazio et al., , 2021). Lately, Orvieto et al., (2022) tackled the issues with unknown $f_{i}^{*}$ and developed a truly adaptive variant. In practice, the SPS method shows excellent empirical performance on overparameterized deep learning models (which satisfy the interpolation condition i.e. $f_{i}^{*}=0,\ \forall i\in[n]$) (Loizou et al., , 2021).
@@ 1.2 Main Contribution
We prove that for smooth non-convex problems with noise having bounded variance KATE has ${\cal O}(\nicefrac{{\log(T)}}{{\sqrt{T}}})$ convergence rate (Theorem 3.4), matching the best-known rates for AdaGrad and Adam (Défossez et al., , 2020).
@@ 2 Motivation and Algorithm Design
Here, we consider the parameter estimation problem in generalized linear models (GLMs) (Nelder and Wedderburn, , 1972; Agresti, , 2015) using maximum likelihood estimation. GLMs are an extension of linear models and encompass several other valuable models, such as logistic (Hosmer Jr et al., , 2013) and Poisson regression (Frome, , 1983), as special cases. The parameter estimation to fit GLM on dataset $\{x_{i},y_{i}\}_{i=1}^{n}$ (where $x_{i}\in\mathbb{R}^{d}$ are feature vectors and $y_{i}$ are response variables) can be reformulated as
@@
for differentiable functions $\varphi_{i}:\mathbb{R}\to\mathbb{R}$ (Shalev-Shwartz and Ben-David, , 2014; Nguyen et al., 2017b, ; Takáč et al., , 2013; He et al., , 2018; Chezhegov et al., , 2024). For example, the linear regression on data $\{x_{i},y_{i}\}_{i=1}^{n}$ is equivalent to solving (4) with $\varphi_{i}(z)=(z-y_{i})^{2}$. Next, the choice of $\varphi_{i}$ for logistic regression is $\varphi_{i}(z)=\log\left(1+\exp{\left(-y_{i}z\right)}\right)$.
@@
respectively, for functions $\varphi_{i}$. In this work, we want to design an algorithm with equivalent performance for the problems (4) and (5). If we can do that, the new algorithm’s performance will not deteriorate for poorly scaled data, i.e., the method will be scale-invariant (Orabona et al., , 2015), which is a special case of affine-invariance, see (Nesterov and Nemirovskii, , 1994; Nesterov, , 2018; d’Aspremont et al., , 2018). To develop such an algorithm, we replace the denominator of AdaGrad step size with its square (remove the square root from the denominator), i.e., $\forall k\in[d]$
$w_{t+1}[k]$ $=$ $w_{t}[k]-\frac{\beta m_{t}[k]}{\sum_{\tau=0}^{t}g_{\tau}^{2}[k]}g_{t}[k]$ (6)
for some $m_{t}\in\mathbb{R}^{d}$. [^2] Sequence $\{m_{t}\}_{t\geq 0}$ can depend on the problem but is assumed to be scale-invariant. The following proposition shows that this method (6) satisfies a scale-invariance property with respect to functional value.
@@
$\hat{w}_{t+1}[k]$ $=$ $\hat{w}_{t}[k]-\frac{\beta m_{t}[k]}{\textstyle{\sum}_{\tau=0}^{t}g_{\tau}^{2}[k]}g_{t}[k],$ (7)
$\quad\hat{w}^{V}_{t+1}[k]$ $=$ $\hat{w}^{V}_{t}[k]-\frac{\beta m_{t}[k]}{\textstyle{\sum}_{\tau=0}^{t}\left(g_{\tau}^{V}[k]\right)^{2}}g^{V}_{t}[k]$ (8)
@@
$\hat{w}_{t}=V\hat{w}^{V}_{t},\quad Vg_{t}=g^{V}_{t},\quad f\left(\hat{w}_{t}\right)=f^{V}\left(\hat{w}^{V}_{t}\right)$ (9)
@@
The Proposition 2.1 highlights that the update rule of the form (6) satisfies a scale-invariance property for GLMs. In contrast, AdaGrad does not satisfy (9) and (10). In Appendix C, we illustrate numerically the scale-invariance of KATE and the lack of the scale-invariance of AdaGrad. We also emphasize that AdaGrad with $\Delta=0$ is known to be a scale-free method [^3] The algorithm is called scale-free if for any $c>0$, it generates the same sequence of points for functions $f$ and $cf$ given the same initialization and hyperparameters. To the best of our knowledge, this definition is introduced by Cesa-Bianchi et al., (2005, 2007) in the context of learning with expert advice and extended to the context of generic online convex optimization by Orabona and Pál, (2015, 2018). We emphasize that scale-freeness and scale-invariance are completely different concepts..
@@
Comparison with the scale-invariant version of AdaGrad by Orabona et al., (2015).
In the special case of GLMs, Orabona et al., (2015) propose a different version of AdaGrad. The method is proposed for the case of online convex optimization, and in the case of standard optimization with GLMs (4), it has the following form
$w_{0}:=0,\quad w_{t+1}:=-\beta\frac{\sum_{\tau=0}^{t}\nabla f_{i_{\tau}}(w_{\tau})}{a_{t}^{2}\sqrt{d}\sqrt{\gamma^{2}+\sum_{\tau=0}^{t}\left(\nicefrac{{\nabla f_{i_{\tau}}(w_{\tau})}}{{a_{\tau}}}\right)^{2}}},\quad a_{t}:=\max\limits_{\tau=0,\ldots,t}|x_{i_{\tau}}|,$ (14)
where $\{i_{\tau}\}_{\tau=0}^{t}$ are arbitrary indices from $[n]$ (e.g., selected uniformly at random), functions $f_{i}:\mathbb{R}^{d}\to\mathbb{R}$ are defined as $f_{i}(w):=\varphi_{i}(x_{i}^{\top}w)$ for $i\in[n]$, and $\gamma$ is such that $f_{i}(w)$ is $\gamma$-Lipschitz for $i\in[n]$. In this setup, the update rule of KATE with $w_{0}=0$ can be written as follows:
$w_{t+1}:=-\beta\sum\limits_{\tau=0}^{t}\frac{m_{\tau}}{b_{\tau}^{2}}\nabla f_{i_{\tau}}(w_{\tau}),\quad m_{t}:=\sqrt{\eta\sum_{\tau=0}^{t}(\nabla f_{i_{\tau}}(w_{\tau}))^{2}+\sum_{\tau=0}^{t}\nicefrac{{(\nabla f_{i_{\tau}}(w_{\tau}))^{2}}}{{b_{\tau}^{2}}}},$
where $b_{t}:=\sqrt{\sum_{\tau=0}^{t}(\nabla f_{i_{\tau}}(w_{\tau}))^{2}}$, $\{i_{\tau}\}_{\tau=0}^{t}$ are sampled from $[n]$ uniformly at random. Although both methods can be seen as variations of AdaGrad due to the terms $\sum_{\tau=0}^{t}\left(\nicefrac{{\nabla f_{i_{\tau}}(w_{\tau}}}{{a_{\tau}}}\right)^{2}$ and $\sum_{\tau=0}^{t}\left(\nabla f_{i_{\tau}}(w_{\tau})\right)^{2}$ respectively, the scale-invariance is achieved quite differently in these methods. The method from (14) uses the feature vectors explicitly in the update rule to ensure scale-invariance: indeed, the square root in the definition of $w_{t+1}$ is independent of scaling, and $a_{t}^{2}$ in the denominator ensures that $\hat{w}_{t+1}=V\hat{w}_{t+1}^{V}$ if we define them similarly to KATE (see equations (7)-(8)). In contrast, KATE achieves the scale-invariance by removing the square root from the denominator (as explained earlier). Moreover, unlike the method from (14), KATE does not use the feature vectors explicitly in its update rule (only in the gradients of $f_{i_{\tau}}$) and, thus, can be used for general stochastic optimization (not necessarily for the case of GLMs).